the sum of two numbers is 2. if one number is subtracted from the other number the result is 8, find the numbers
Answer by ankor@dixie-net.com(22740)

the sum of two numbers is 2.
a + b = 2
if one number is subtracted from the other number the result is 8,
a - b = 8
:
find the numbers
Use elimination here
a + b = 2
a - b = 8
--------------Addition eliminates b, find a
2a = 10
a = 10/2
a = 5 is one number
:
Find b, replace a with 5 in the first equation
5 + b = 2
b = 2 - 5
b = -3 is the other number
;
:
Check solutions in the 2nd equation
5 - (-3) = 8